OpenStudy (anonymous):

I need help with recipe conversion...need to convert a 12 serving recipe to 50 servings. Those who help get a medal, a new fan, and virtual mini cheesecakes! :D

OpenStudy (anonymous):

Mini Cheesecakes Yield: 12 servings 12 vanilla wafers 2 (8 ounce) packages cream cheese, softened 2 eggs 2 tablespoons lemon juice 2/3 cup white sugar 1 (21 ounce) can cherry pie filling

OpenStudy (anonymous):

Need to convert the above recipe so that it makes 50 servings. How would I do this?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Hello @texaschic101! Can you help? :)

OpenStudy (texaschic101):

12/12 = x/50 (12 vanilla wafers to 12 serv. = x wafers to 50 servings) cross multiply 12x = (50)(12) 12x = 600 x = 600/12 x = 50 you will need 50 vanilla wafers 2/12 = x/50 (2 packs cream cheese to 12 serv. = x packs to 50 servings) cross multiply 12x = (50)(2) 12x = 100 x = 100/12 x = 8.33 or 8 1/3 packs you will need 8.33 packs of cream cheese you see how I am setting these up ? Do you understand ?

OpenStudy (anonymous):

I see, thank you, I was having some difficulty setting up my proportions. Here you go...50 virtual mini cheesecakes. :D

OpenStudy (texaschic101):

yummy :)

OpenStudy (anonymous):

lol

OpenStudy (texaschic101):

if you have any more problems with this, just tag me :)